# Revisiting Einstein’s analogy: black holes as gradient-index lenses

### v1

@@ II Photon trajectories around equatorial Kerr black holes
The resulting trajectories play the same role as the radial geodesics in the Schwarzschild geometry Chandrasekhar_1998 . Fig. 1 shows radial geodesics for a Kerr black hole rotating clockwise for $a=0.1m_{b}$, and $0.9m_{b}$.
Figure 1: Radial geodesics $(h=a)$ are shown for the cases $(a)$ $a=0.1$ and $(b)$ $a=0.9$ , using Planck units and $m_{b}=1$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
According to Ref. 29, Eq. (9) can be integrated exactly in terms of elliptic integrals of the first and third kinds gradshteyn2000table when integrating from the point $r_{bl,0}$, where the distance from the geodesic to the event horizon is minimized, to the position $r_{bl}$, with $r_{bl}>r_{bl,0}$. By reparametrizing $r_{bl}=2m_{b}x$, $a=2m_{b}j$, and $h=2m_{b}\psi$, where $x,j,\psi$ are dimensionless, this condition of minima distance is equivalent to fan2018analytical
@@
where $\pm$ indicates prograde and retrograde trajectories. The explicit form of the exact solution to these null geodesics can be found in Ref. 29. However, it is important to note a misprint in Eq. (17) in that article: the variable $z$ should be replaced with $\sin(z)$. Under these conditions, it is straightforward to demonstrate that when $j=0$, the results in Ref. 29 reduce to the Schwarzschild case, as shown in Ref. 35. Fig. 2 presents the analytical prograde and retrograde geodesics for the cases $(a)$ $a=0.1m_{b}$ and $(b)$ $a=0.9m_{b}$ with $r_{bl,0}=4m_{b}$.
@@
Figure 2: Exact prograde and retrograde paths and numerical null geodesics for the cases $(a)$ $a=0.9$ and $(b)$ $0.99$, using Planck units and $m_{b}=1$, and the point of convergence at $r_{bl,0}=4$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
As we are interested in reproducing light trajectories around a black hole using a symmetric GRIN media, we mainly focus on trajectories that cross the event horizon. To obtain such trajectories through numerical integration, we first need to ensure that the coefficient of $dr$ in Eq. (9) is a real-valued function, which implies that the values of $\psi$ must be restricted. This restriction arises from the condition $f(x)>0$ for the interval $x\in(0,\infty)$, which can be reduced to
@@
Thus, by solving for $\psi$, it is found the range of vaues that ensures the numerical integration of Eq. (9) is well-defined:
@@
[…]
@@
[…]
Here, $x_{0}$ is chosen to be arbitrarily close to the exterior event horizon. In Fig. 3, it is shown the numerically computed paths using the trapezoidal method for $a=0m_{b}$ and $0.9m_{b}$ by taking different values of $\psi$ in the range Eq. (17); all trajectories converge at $r_{bl_{f}}=9.5m_{b}$.
Figure 3: Numerical null geodesics for the cases $(a)$ $a=0$ and $(b)$ $0.9$, using Planck units and $m_{b}=1$, and the position of convergence at $r_{bl,f}=9.5$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
@@ III Fermat’s ray invariant for GRIN media with spherical symmetry
Fermat’s principle can be expressed using orthogonal generalized coordinates associated with an arbitrary curvilinear orthogonal coordinate system that depends on a parameter $\tau$ $(q_{1}(\tau),q_{2}(\tau),q_{3}(\tau))$ as follows:
$\delta\int_{C}\mathcal{L}(q_{1},q_{2},q_{3};\dot{q_{1}},\dot{q_{2}},\dot{q_{3}},\tau)\text{d}\tau=0,$
where, $\mathcal{L}(q_{1},q_{2},q_{3};\dot{q_{1}},\dot{q_{2}},\dot{q_{3}},\tau)$ is the optical Lagrangian, $\dot{q}_{i}=dq_{i}/d\tau$, with $i=1,2,3$, and $C$ represents the light-ray trajectories parameterized by $\tau$.
@@
When using a spherical coordinate system $(r,\theta,\phi)$ and a radially symmetric GRIN medium, it is sufficient to analyze the behavior of light along the equator $(\theta=\pi/2)$, since any trajectories lie in a plane passing through the origin. This is the primary reason why the equatorial null geodesics of a Kerr black hole will be modeled using a spherical GRIN medium. Under these conditions, and by using $r$ as the parameter of the trajectories, the Lagrangian function is given by:
@@
with $\phi_{r}=d\phi/dr$. Thus, the light-ray trajectories satisfy the Euler-Lagrange equations Lakshminarayanan , and since the Lagrangian does not explicitly depend on $\phi$, there is a conserved quantity given by Gomez-Correa:21 ; G_mez_Correa_2023 :
$K=rn(r)\sin{\varphi},$ (23)
where $\varphi$ is the angle between the vector $\hat{{\bf s}}$, tangent to the ray, and $\bf{r}$, as illustrated in Fig. 4. In optics, Eq. (23) is referred to as Fermat’s ray invariant for inhomogeneous media with spherical symmetry Gomez-Correa:21 ; Gomez-Correa:22 or as Bouguer’s formula Born_2019 .
It is important to emphasize that the value of $K$, once defined, does not change as the ray of light propagates. In other words, $K$ is a conserved quantity along the light ray.
Equation (23) shows that Fermat’s ray invariant depends on variations in the GRIN distribution. Therefore, the GRIN distribution can be determined if the value of $K$ is known and the trajectory is given (thus, $\varphi$ can be determined), as follows:
$n(r)=\frac{K}{r\sin{\varphi}},$ (24)
The expression above forms the core of the PhysGRIN method for spherical symmetry. The numerical implementation to determine the refractive index for photon trajectories around a Kerr black hole will be analyzed in Sec. V.
@@ IV Analytical GRIN reconstruction from photon trajectories
To determine the value of $K$, it is essential to know $\sin{\varphi}$. Consequently, as $\varphi$ is the angle between $\hat{\bf s}$ and $\bf{r}$, the following relationship holds:
$\sin\varphi=\frac{r(\phi)}{\sqrt{r^{2}(\phi)+\left(\frac{dr}{d\phi}\right)^{2}}}.$ (25)
@@
$n(r)=\frac{K\sqrt{r^{2}+(dr/d\phi)^{2}}}{r^{2}}.$ (26)
Now, let us assume that $n_{e}=n(R)$ and that $dr/d\phi$ can be expressed as a function of $r$, that is, $dr/d\phi=q(r)$. From Eq. (26), at $r=R$, we can derive $K$:
$K=\frac{R^{2}n_{e}}{\sqrt{R^{2}+q^{2}(R)}}.$ (27)
@@
If we now consider the Boyer-Lindquist coordinates, it follows that:
$q(r)=\frac{r_{bl}}{\sqrt{r_{bl}^{2}+a^{2}}}\frac{dr_{bl}}{d\phi}.$ (29)
Using Eq. (9), we find that for the most general case, the index of refraction is given by:
@@
$s(r)$ $=$ $\left[r^{2}-2m_{b}\sqrt{r^{2}-a^{2}}\right]^{2}$ (31)
×$\times\left[(r^{2}-a^{2})(r^{2}-h^{2})+2m_{b}\sqrt{r^{2}-a^{2}}(a-h)^{2}\right],$
@@
Equation (30) establishes the analytical distribution of the refractive index when the trajectories are represented in Cartesian coordinates $x_{c}$, $y_{c}$. Thus, Eq. (30) agrees with the observation made in Ref. 18 that the most accurate definition of the refractive index is achieved when light rays in optical media coincide with null geodesics.
@@
Although the same approach to obtain the analytical form of the index of refraction was reported in Ref. 14, our results differ from theirs because they represented the null geodesics in Boyer-Lindquist coordinates (that is to say, the GRIN medium depends on the coordinate system used to visualize the trajectories); Additionally, they do not provide an explicit form of the proportionality constant $K$.
We use Cartesian coordinates instead of Boyer-Lindquist coordinates because, in such a representation, the singularity of the Kerr black hole appears as an annulus, which is a common depiction.
From Eq. (28), it is clear that the radial distribution of the GRIN medium depends on the nature of the function $q(r)$. This function’s existence and explicit form depend on the origin of the ray’s trajectory. Determining this function can be particularly challenging if we focus on the paths of photons around equatorial black holes. This task requires prior knowledge of the theory of relativity and tensor calculus, as well as an understanding of the nature and geometric properties of spacetime, as discussed in Sec. II.
@@ V Reconstruction without general relativity
So far, in Sec. IV, we have outlined the procedure to obtain the analytical form of the GRIN medium that reproduces null geodesics around equatorial Kerr black holes. Obtaining $q(r)$ in a broader context requires understanding general relativity, which can be challenging for a non-expert. Moreover, the analysis can take significant time, which may not be necessary if the main goal is reconstructing the GRIN medium. We will now explain how using Eq. (24) enables a remarkable simplification of the process when focusing exclusively on the trajectory.
Recent findings demonstrate that the value of $K$ can be determined independently of the GRIN distribution, as shown below Gomez-Correa:22 :
$K=Rn_{e}\sin{\theta_{Inc}},$ (33)
where $n_{e}$ is the external refractive index in which the GRIN distribution is immersed, and $\theta_{Inc}$ is the angle of incidence of the straight light ray over the GRIN surface at the position ${\bf r}_{0}$, as shown in Fig. 4.
@@
We assume the ray is discretized to develop a systematic method for reconstructing the GRIN. Consider a discretized ray passing through a spherical GRIN lens, as illustrated in Fig. 4. Let ${\bf r}_{i}=(x_{i},y_{i})$ represent the points along the ray path, where $i$ denotes the discretization index. According to Eq. (34), the value of the index of refraction at point ${\bf r}_{i}$ along the ray’s trajectory is given by
@@
where $r_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}}$. This expression can be quickly evaluated when the trajectory is known because $\varphi(r_{i})$ is calculated as G_mez_Correa_2023
@@
In Sec. VI, the PhysGRIN method is applied to reconstruct the GRIN medium representing the equatorial spacetime of a Kerr black hole. These reconstructions are then compared to the analytical solution provided by Eq. (30).

### v2

@@ II Photon trajectories in the equatorial plane of Kerr black holes
The resulting trajectories, characterized by the vanishing of the Carter constant $K$ Carter_1968 (i.e., $K=\varepsilon^{2}(a-h)^{2}$ at $\theta=\pi/2$ and $\dot{\theta}=0$, which represents a conserved quantity that arises from the unexpected separability of the Hamilton-Jacobi equation), fulfill a similar role to that of radial geodesics in Schwarzschild geometry Chandrasekhar1983mathematical . Fig. 1 shows such geodesics for a Kerr black hole rotating clockwise for $a=0.1m_{b}$, and $0.9m_{b}$.
Figure 1: Geodesics with vanishing Carter constant are shown for the cases $(a)$ $a=0.1$ and $(b)$ $a=0.9$, using Planck units and $m_{b}=1$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
Equation (9) can be solved exactly using elliptic integrals, as first shown in the context of geodesic motion in the Schwarzschild metric by Hagihara hagihara1930theory and later discussed in Chandrasekhar’s classical treatment Chandrasekhar1983mathematical . Specifically, Equation (9) can be integrated exactly in terms of elliptic integrals of the first and third kinds fan2018analytical ; gradshteyn2000table when integrating from the point $r_{bl,0}$, where the distance from the geodesic to the event horizon is minimized, to the position $r_{bl}$, with $r_{bl}>r_{bl,0}$. By reparametrizing $r_{bl}=2m_{b}x$, $a=2m_{b}j$, and $h=2m_{b}\psi$, where $x,j,\psi$ are dimensionless, this condition of minima distance is equivalent to fan2018analytical
@@
where $\pm$ indicates prograde (+) and retrograde (-) trajectories. The explicit form of the exact solution to these null geodesics can be found in Ref. fan2018analytical . However, it is important to note a misprint in Eq. (17) in that article: the variable $z$ should be replaced with $\sin(z)$. Under these conditions, it is straightforward to demonstrate that when $j=0$, the results in Ref. fan2018analytical reduce to the Schwarzschild case, as shown in Ref. Bret_n_2017 . Fig. 2 presents the analytical prograde and retrograde geodesics for the cases $(a)$ $a=0.1m_{b}$ and $(b)$ $a=0.9m_{b}$ with $r_{bl,0}=4m_{b}$.
@@
Figure 2: Exact prograde and retrograde paths and numerically calculated null geodesics for the cases $(a)$ $a=0.9$ and $(b)$ $0.99$, using Planck units and $m_{b}=1$, and the point of convergence at $r_{bl,0}=4$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
As we are interested in reproducing light trajectories around a black hole using a symmetric GRIN medium, we will focus on null geodesics that approach the event horizon and experience strong deflection. In Boyer–Lindquist coordinates, light rays do not cross the event horizon in finite coordinate time; instead, they asymptotically approach it. This coordinate behavior does not affect the physical equivalence of the modeled trajectories in the GRIN medium. To obtain such trajectories through numerical integration, we first need to ensure that the coefficient of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r$ in Eq. (9) is a real-valued function, which implies that the values of $\psi$ must be restricted. This restriction arises from the condition $f(x)>0$ for the interval $x\in(0,\infty)$, which can be reduced to
@@
Thus, by solving for $\psi$, it is found that the range of values ensuring the well-defined numerical integration of Eq. (9) is given by:
@@
$\phi(x)=\int_{x_{0}}^{x}\frac{x^{\prime}(\psi x^{\prime}+j-\psi)}{(x^{\prime 2}+j^{2}-x^{\prime})\sqrt{f(x^{\prime})}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}x^{\prime}-\phi_{0},$ (20)
@@
$\phi_{0}=\int_{x_{0}}^{x_{f}}\frac{x^{\prime}(\psi x^{\prime}+j-\psi)}{(x^{\prime 2}+j^{2}-x^{\prime})\sqrt{f(x^{\prime})}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}x^{\prime}.$ (21)
Here, $x_{0}$ is chosen to be arbitrarily close to the exterior event horizon. In Fig. 3, it is shown the numerically computed paths using the trapezoidal method for $a=0m_{b}$ and $0.9m_{b}$ by taking different values of $\psi$ in the range $(\psi_{1}(j),\psi_{2}(j))$; all trajectories converge at $r_{bl_{f}}=9.5m_{b}$.
Figure 3: Numerically calculated null geodesics for the cases $(a)$ $a=0$ and $(b)$ $0.9$, using Planck units and $m_{b}=1$, and the position of convergence at $r_{bl,f}=9.5$. Cartesian coordinates $x_{c}$ and $y_{c}$ were used.
@@ III Fermat’s ray invariant for GRIN media with spherical symmetry
Fermat’s principle can be expressed using orthogonal generalized coordinates associated with an arbitrary curvilinear orthogonal coordinate system that depends on a parameter $\eta$ $(q_{1}(\eta),q_{2}(\eta),q_{3}(\eta))$ as follows:
$\delta\int_{C}\mathcal{L}(q_{1},q_{2},q_{3};\dot{q_{1}},\dot{q_{2}},\dot{q_{3}},\eta)\text{d}\eta=0,$
where, $\mathcal{L}(q_{1},q_{2},q_{3};\dot{q_{1}},\dot{q_{2}},\dot{q_{3}},\eta)$ is the optical Lagrangian, $\dot{q}_{i}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}}q_{i}/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\eta$, with $i=1,2,3$, and $C$ represents the light-ray trajectories parameterized by $\eta$.
@@
When using a spherical coordinate system $(r,\theta,\phi)$ and a radially symmetric GRIN medium, it is sufficient to analyze the behavior of light along the equator $(\theta=\pi/2)$, since any trajectories lie in a plane passing through the origin. This is the primary reason why the equatorial null geodesics of a Kerr black hole will be modeled using a spherical GRIN medium thus, only a thin GRIN disk, where the light rays lie, is required Tinguely_2020 . Under these conditions, and by using $r$ as the parameter of the trajectories, the Lagrangian function is given by:
@@
with $\phi_{r}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r$. Thus, the light-ray trajectories satisfy the Euler-Lagrange equations, and since the Lagrangian does not explicitly depend on $\phi$, there is a conserved quantity, $K_{r}\equiv\partial\mathcal{L}/\partial\phi_{r}$, known as Fermat’s ray invariant, given explicitly by Gomez-Correa:21 ; Lakshminarayanan ; G_mez_Correa_2023 :
$K_{r}=rn(r)\sin{\varphi},$ (23)
where $\varphi$ is the angle between the vector $\hat{{\bf s}}$, tangent to the ray, and $\bf{r}$, as illustrated in Fig. 4. In optics, Eq. (23) is referred to as generalized Snell’s law for inhomogeneous media with spherical symmetry Luneburg_1964 ; Lakshminarayanan ; Gomez-Correa:21 ; Gomez-Correa:22 or as Bouguer’s formula Born_2019 . The value of $K_{r}$ is determined at the entrance point of the ray into the GRIN medium, at $r=R$. It is important to emphasize that since $K_{r}$ is an invariant or conserved quantity, once defined it does not change as the light ray propagates.
Eq. (23) shows that Fermat’s ray invariant depends on variations in the GRIN distribution. Therefore, the GRIN distribution can be determined if the value of $K_{r}$ is known and the trajectory is given (thus, $\varphi$ can be determined), as follows:
$n(r)=\frac{K_{r}}{r\sin{\varphi}}.$ (24)
The expression above forms the core of the PhysGRIN method for spherical symmetry. The numerical implementation to determine the refractive index for photon trajectories around a Kerr black hole will be discussed in Sec. V.
@@ IV Analytical GRIN reconstruction from photon trajectories
To determine the value of $K_{r}$, it is essential to know $\sin{\varphi}$. Consequently, as $\varphi$ is the angle between $\hat{\bf s}$ and $\bf{r}$, the following relationship holds:
$\sin\varphi=\frac{r(\phi)}{\sqrt{r^{2}(\phi)+\left(\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi}\right)^{2}}}.$ (25)
@@
$n(r)=\frac{K_{r}\sqrt{r^{2}+({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi)^{2}}}{r^{2}}.$ (26)
Now, let us assume that $n(R)=n_{e}$, and that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi$ can be expressed as a function of $r$, that is, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r/{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi=q(r)$. From Eq. (26), at $r=R$, we can derive $K_{r}$:
$K_{r}=\frac{R^{2}n_{e}}{\sqrt{R^{2}+q^{2}(R)}}.$ (27)
@@
Now, if we consider the Boyer-Lindquist coordinates in the equatorial plane ($\theta=\pi/2$), their relation to spherical coordinates is given by $r=\sqrt{r_{bl}^{2}+a^{2}}$. Thus, we obtain:
$q(r)=\frac{r_{bl}}{\sqrt{r_{bl}^{2}+a^{2}}}\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}r_{bl}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\text{d}}\phi}.$ (29)
By substituting Eq. (9) into the last equation and Eq. (29) into Eq. (28), and then expressing the index of refraction in spherical coordinates, where $r_{bl}=\sqrt{r^{2}-a^{2}}$, we find that, in the most general case, the index of refraction is given by:
@@
$s(r)$ $=$ $\left[r^{2}-2m_{b}\sqrt{r^{2}-a^{2}}\right]^{2}\left[(r^{2}-a^{2})(r^{2}-h^{2})+2m_{b}\sqrt{r^{2}-a^{2}}(a-h)^{2}\right],$ (31)
@@
Equation (30) establishes the analytical distribution of the refractive index when the trajectories are represented in Cartesian coordinates $x_{c}$, $y_{c}$. Thus, Eq. (30) agrees with the observation made in Ref. Ramezani_Aval_2024 that the most accurate definition of the refractive index is achieved when light rays in optical media coincide with null geodesics.
@@
Although the same approach to obtaining the analytical form of the index of refraction was reported in Ref. Tinguely_2020 , they do not provide an explicit form of the proportionality constant $K_{r}$. Moreover, the main difference of our procedure from theirs is that they represented the null geodesics in Boyer-Lindquist coordinates, thus avoiding the change of variables in Eq. (29), indicating that spherical coordinates and Boyer-Lindquist coordinates are not related by a conformal transformation (i.e., the GRIN medium depends on the coordinate system used to describe the trajectories). An alternative approach to obtaining the index of refraction that exactly reproduces photon trajectories for static spherically symmetric spacetimes involves introducing isotropic coordinates, where the curved spacetime becomes conformally flat Nouri_Zonoz_2022 . However, this approach is not generally applicable for stationary spacetimes Parvizi_2024 ; therefore, the approach presented here and the numerical method discussed in the next section offers a more general approach to reconstruct GRIN media. Here, we decided to use Cartesian coordinates instead of Boyer-Lindquist coordinates because, in such a representation, the singularity of the Kerr black hole appears as an annulus, which is a common depiction.
From Eq. (28), it is clear that the radial distribution of the GRIN medium depends on the nature of the function $q(r)$. The function’s existence and explicit form depend on the origin of the ray’s trajectory. Determining this function can be particularly challenging if we focus on the paths of photons lying in the equatorial plane of a black hole. This task requires prior knowledge of the theory of relativity and tensor calculus, as well as an understanding of the nature and geometric properties of spacetime, as discussed in Sec. II.
@@ V Reconstruction without general relativity
So far, in Sec. IV, we have outlined the procedure to obtain the analytical form of the GRIN medium that reproduces null geodesics lying in the equatorial plane of a black hole. Obtaining $q(r)$ in a broader context requires understanding GR, which can be challenging for a non-expert. Moreover, the analysis can take significant time, which may not be necessary if the main goal is reconstructing the GRIN medium. We will now explain how using Eq. (24) enables a remarkable process simplification when focusing the attention exclusively on the trajectory.
Recent findings demonstrate that the value of $K_{r}$ can be determined independently of the GRIN distribution, as shown below Gomez-Correa:22 :
$K_{r}=Rn_{e}\sin{\theta_{Inc}},$ (33)
where $n_{e}$ is the external refractive index in which the GRIN distribution is immersed, and $\theta_{Inc}$ is the angle of incidence of the straight light ray over the GRIN surface at the position ${\bf r}_{0}$ ($|{\bf r}_{0}|=R$), as shown in Fig. 4.
@@
We assume the ray is discretized to develop a systematic method for reconstructing the GRIN. Consider a discretized ray passing through a spherical GRIN lens, as illustrated in Fig. 4. Let ${\bf r}_{i}=(x_{i},y_{i})$ represent the points along the ray path, where $i$ denotes the discretization index. According to Eq. (34), the value of the index of refraction at the point ${\bf r}_{i}$ along the ray’s trajectory is given by
@@
where $r_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}}$. This expression can be quickly evaluated when the trajectory is known because $\varphi(r_{i})$ can be calculated from G_mez_Correa_2023
@@
In Sec. VI, the PhysGRIN method is applied to reconstruct the GRIN medium representing an equatorial slice of the spacetime of a Kerr black hole. These reconstructions are then compared to the analytical solution provided by Eq. (30).